OpenStudy (anonymous):

You are hiring someone to prune your trees. One charges $15 consult fee + 8/hr. Another charges no consult fee and $10/hr. who would you hire for a 3 hr job? where do they break even?

OpenStudy (anonymous):

1) 3 hours work: 1st consultant: 15+3*8=39 2nd consultant: 3*10=30

OpenStudy (anonymous):

2nd consultant is cheaper right?

OpenStudy (anonymous):

y=10(x) y=8x+15

OpenStudy (anonymous):

this would work to calculate where do they break even. 8x+15=10x x=... hours

OpenStudy (anonymous):

they break even at 10x=8x+15 ... 2x=15 .... x=7.5

OpenStudy (anonymous):

yep. 7.5 hours